# The stellar bar – dark matter halo connection in the TNG50 simulations

### v2

@@ Abstract
Stellar bars in disk galaxies grow as stars in near circular orbits lose angular momentum to their environments, including their Dark Matter (DM) halo, and transform into elongated bar orbits. This angular momentum exchange during galaxy evolution hints at a connection between bar properties and the DM halo spin $\lambda$, the dimensionless form of DM angular momentum. We investigate the connection between halo spin $\lambda$ and galaxy properties in the presence/absence of stellar bars, using the cosmological magneto-hydrodynamic TNG50 simulations at multiple redshifts ($10<z_{r}<$ 1). We determine the bar strength (or bar amplitude, $A_{2}/A_{0}$), using Fourier decomposition of the face-on stellar density distribution. We determine the halo spin for barred and unbarred galaxies ($0.70<A_{2}/A_{0}<$ 0.7) in the centre of the DM halo, close to the galaxy stellar disk. At $z_{r}=0$, there is an anti-correlation between halo spin and bar strength. Strongly barred galaxies ($A_{2}/A_{0}>0.4$) reside in DM halos with low spin and low specific angular momentum at their centers. In contrast, unbarred/weakly barred galaxies ($A_{2}/A_{0}<0.2$) exist in halos with higher central spin and higher specific angular momentum. The anti-correlation is due to the barred galaxies’ higher DM mass and lower angular momentum than the unbarred galaxies at $z_{r}=0$, as a result of galaxy evolution. At high redshifts ($z_{r}=1$), all galaxies have higher halo spin compared to those at lower redshifts ($z_{r}=0$), with a weak anti-correlation for galaxies having $A_{2}/A_{0}>0.2$. The formation of a DM bar in strongly barred systems highlights how angular momentum transfer to the halo can influence its central spin.
@@ 1 Introduction
Investigating the properties of halo spin and its connection with galaxy properties has been of interest in many studies (Bett et al., 2007, 2010; Bett & Frenk, 2012, 2016). Halo spin plays an important role determining galaxy sizes, as faster-rotating halos having higher halo spins are associated with larger stellar disks (Kim & Lee, 2013; Jiang et al., 2019; Rodriguez-Gomez et al., 2022; Pérez-Montaño et al., 2022; Yang et al., 2023). Romeo et al. (2023) studied a large sample of barred S0 galaxies and found a connection between halo spin, disk spin and bar structure. Zjupa & Springel (2017) found that the halo spin (both $\lambda_{B}$ and $\lambda_{P}$ at $r_{vir}$) are not much affected by the collapse of the baryons in the DM halos. Recently, Rosas-Guevara et al. (2022) using current cosmological hydrodynamical TNG50 simulations (Nelson et al., 2019; Pillepich et al., 2019) showed that the halo spin (at $r_{vir}$; $\lambda_{B}$) of barred galaxies is lower than that of the unbarred galaxies at redshift $z_{r}=0$ (also see Izquierdo-Villalba et al. 2022). However, there has been no comprehensive study on the origin of the connection between bar strength and halo spin or halo angular momentum. We will see later in Section 4.1, that the angular momentum of the stellar disk is more than the angular momentum of the DM halo in the central region of the galaxies (for example, 10 times more inside $r<10$ kpc). Only a significant transfer in angular momentum from the disk to the DM halo can lead to a halo spin change in the halo’s inner region. However, no amount of angular momentum transfer can change the halo spin within the virial radius of the halo.
@@ 2 TNG50 Simulations and galaxy sample
We avoid galaxies that do not have enough star particles to resolve the central region of the disk. Following the above criteria, we select barred and unbarred disk galaxies from the above samples, with 597 galaxies at redshift $z_{r}=0$, 650 galaxies at $z_{r}=0.1$ and 509 galaxies at $z_{r}=1.0$. Note that we do not track the progenitors of the galaxies at higher redshifts but construct independent samples of galaxies at the three redshifts.
@@ 2.1 Barred and unbarred galaxies
Figure 2: The anti-correlation between bar strength and dark matter halo properties at redshift $z_{r}=0.0$. Figure shows the median halo spin (panel a), median DM angular momentum (panel b), median DM mass (panel c) and median specific angular momentum of DM (panel d) for each of the bar strength groups from Table 1 at redshift $z_{r}=0$, for two radii, $r_{m}=5$ kpc (green, Sample 1r5), 10 kpc (magenta, Sample 1r10), in the velocity range $200140<\rm Max\left[V_{c,DM}(r<r_{m})\right]/(kms^{-1})<$ 200. The solid circles denote the median values and the shaded regions are bounded by the $16^{th}$ and $84^{th}$ percentile curves.
@@ 3 Halo spin of Barred and Unbarred galaxies
We aim to study the halo spin of barred and unbarred galaxies having similar masses of their DM halos in the central region and close to the baryonic disk. We use the circular velocity ${\rm V_{c,DM}}(r)=\sqrt{GM_{DM}(r)/r}$ of the DM component to select galaxies with similar DM halo mass profiles ($M_{DM}(r)$). We use the selection criteria of $200140<\rm Max\left[V_{c,DM}(r<r_{m})\right]/(kms^{-1})<$ 200 at radii $r_{m}=5$ and 10 kpc, on each of the seven bar strength groups at $z_{r}=0$ from Table 1, as it maximises the number of galaxies following this selection criterion. At $z_{r}=0$, we find 25, 21, 42, 34, 29, 29, and 10 galaxies at $r_{m}=10$ kpc (Sample 1r10, hereafter) and 19, 22, 48, 44, 45, 33 and 9 galaxies at $r_{m}=5$ kpc (Sample 1r5, hereafter) for each of the bar strength intervals from Table 1. Except for the highest bar strength bin ($0.70.6<A_{2}/A_{0}<$ 0.7), all other bar strength intervals have a moderate number of galaxies.
@@ 3.1 Galaxies with similar DM circular velocities at redshift $z_{r}=0$
In Figure 2, we present the median halo spin $\lambda(r)$ (panel a), median DM angular momentum $L_{DM}(r)=\sqrt{L^{2}_{x}+L^{2}_{y}+L^{2}_{z}}$ (panel b), median DM mass $M_{DM}(r)$ (panel c) and the median DM specific angular momentum $L_{DM}(r)/M_{DM}(r)$ (panel d) using Sample 1r5 (green) and Sample 1r10 (magenta), with their median bar strength in the x-axis of each panel. The solid circles show the median values of each of the above quantities, and shaded regions are bounded by $16^{th}$ and $84^{th}$ percentile curves.
@@
Figure 3: Halo spin of strongly barred galaxies ($A_{2}/A_{0}>0.3$) consistently decreases from high to low redshifts, $z_{r}=1$ to 0. Figure shows the evolution of halo spin $\lambda$ (panel a), DM halo mass $M_{DM}$ (panel b), DM halo angular momentum $L_{DM}$ (panel c), DM specific angular momentum $L_{DM}/M_{DM}$ (panel d), stellar to DM mass ratio $M_{\star,90}/M_{DM,90}$ within $R_{\star,90}$ (panel e), stellar mass $M_{\star,90}$ (panel f), stellar angular momentum $L_{\star,90}$ (panel g) and stellar disk specific angular momentum $L_{\star,90}/M_{\star,90}$ (panel h) of the galaxies in Sample 1r10 at $z_{r}=0$ (magenta), 0.2 (green), 0.5 (blue), 0.7 (orange) and 1 (grey). To avoid crowding we show three redshifts in all panels except in panel a.
@@ 4.2 Correlations between bar strength and stellar and DM angular momentum
Figure 6: Increase in bar strength leads to the decrease in disk angular momentum and coincides with the rise in DM angular momentum. X-axis of all panels: (a) Pearson correlation coefficient $R$ for stellar angular momentum $L_{\star}$ and bar strength $A_{2}/A_{0}$; (b) DM angular momentum $L_{DM}$ and $L_{\star}$; (c) $L_{\star}$ and stellar mass $M_{\star}$; (d) $L_{DM}$ and DM mass $M_{DM}$ and (e) $L_{DM}$ and $A_{2}/A_{0}$. All quantities are estimated within $r=10$ kpc. The y-axes show the number of systems for barred (blue) and unbarred galaxies (yellow).
@@ 4.3 Shadow Dark Matter Bar
[…]
@@ 5 Halo spin of galaxies in Sample-2 and Sample-3
We investigate if there is an anti-correlation between the halo spin and the bar strength for galaxies with comparable DM masses at high redshifts, similar to $z_{r}=0$ (Section 2.1). Here we examine the barred and unbarred galaxies at higher redshifts $z_{r}=0.1$ and 1.0 by forming samples of galaxies having similar DM mass using Sample 2 and Sample 3 from Table 1. We apply the same selection criteria from Section 3.1 based on the DM circular velocities: $200140<\rm Max\left[V_{c,DM}(r<r_{m})\right]/(kms^{-1})<$ 200, on each of the seven bar strength groups in Sample 2 at $z_{r}=0.1$ and Sample 3 at $z_{r}=1$. We find 37, 58, 46, 39, 40, 28 and 12 galaxies at $r_{m}=10$ kpc (Sample 2r10 hereafter) and 37, 40, 42, 42, 49, 34, 12, galaxies at $r_{m}=5$ kpc (Sample 2r5 hereafter). At $z_{r}=1.0$, we find lesser number of galaxies in Sample-3 in the above circular velocity range: 5, 7, 11, 19, 23, 28 and 11 galaxies for $r_{m}=5$ kpc (Sample 3r5 hereafter) and 8, 19, 15, 24, 33, 38 and 15 galaxies for $r_{m}=10$ kpc (Sample 3r10 hereafter). We present the median halo spin, median DM angular momentum, median DM mass and median DM specific angular momentum for the seven bar strength bins in Figure 8, with Samples 2r5 (panels a–d, green), 2r10 (panels a–d, magenta), 3r5 (panels e–h, green) and 3r10 (panels e–h, magenta), (similar to Figure 2). We have also conducted tests with a slightly different circular velocity selection criterion of $180130<\rm Max\left[V_{c,DM}(r<r_{m})\right]/(kms^{-1})<$ 180 with larger number of galaxies at $z_{r}=1$ to find 18, 24, 16, 39, 33, 32 and 14 galaxies for $r_{m}=10$ kpc and 8, 12, 12, 24, 30, 35 and 15 galaxies for $r_{m}=5$ kpc for the seven bar strength bins and found similar trends in the results.
@@ 6.1 Sample selection and biases
Figure 9: Dark matter halo mass (left column), angular momentum (middle column) and energy (right column) of the sample galaxies at redshifts $z_{r}=$0.0 (top row), 0.1 (middle row) and 1.0 (bottom row), within spherical radius 10 kpc. In each panel, grey points are values of each quantity divided into seven bar strength groups (Table 1). Blue, green and purple lines trace the median values in each bar strength group, and the shaded regions show the 16th and 84th percentile of the sample. Each bar strength interval consists of DM halos with a wide range of mass, angular momentum and energy, however, the scatter decreases at high redshifts.
@@ 6.2 Convergence
The galaxies that do not satisfy the convergence criteria at $r_{m}=R_{\star,60}$ are mostly the low-mass galaxies that are not included in the galaxy samples in Sections 3.1 and 5, where we use $r_{m}=5$ and 10 kpc and use massive galaxies with $V_{c,DM}(r<r_{m})>140$ km s-1. Low-mass galaxies have small sizes. For low-mass galaxies, $R_{\star,60}$ are smaller compared to the high-mass galaxies (see Figure 13 in Appendix C). We exclude these low-mass galaxies while estimating halos spin for $r_{m}=R_{\star,60}$ at $z_{r}=0.0,$ 0.1 and 1.0 in Figure 14 (panels c, f and i ) in Appendix D. For other values of $r_{m}$ (i.e., 5 kpc, 10 kpc, $R_{\star,90}$, $R_{200}$, $0.15R_{200}$ and $0.05R_{200}$) the low-mass galaxies satisfy the convergence criteria. For more details on how we avoid the effects of numerical relaxation at small radii in our samples, see Appendix C.
@@ Appendix A Halo spin in massive and low-mass galaxies
We study the DM halo spin for comparable DM mass galaxies within the velocity range of $250200<\rm Max\left[V_{c,DM}(r<r_{m})\right]/(kms^{-1})<$ 250 for $r_{m}=5$ and 10 kpc. We find 0, 6, 21, 23, 33, 8 and 0 number of galaxies for $r_{m}=10$ kpc for the seven bar strength ranges in Table 1 and similarly for $r_{m}=5$ kpc we have 0, 10, 19, 19, 17, 3 and 0 galaxies for each of the seven bar strength intervals. We avoid low statistics by considering the bar strength intervals with $\geq 6$ galaxies. In Figure 10 we present the halo spin (panel a), halo angular momentum (panel b), halo mass (panel c) and halo specific angular momentum (panel d) for the above sample of barred and unbarred galaxies.

### v3

@@ Abstract
Stellar bars in disk galaxies grow as stars in near circular orbits lose angular momentum to their environments, including their Dark Matter (DM) halo, and transform into elongated bar orbits. This angular momentum exchange during galaxy evolution hints at a connection between bar properties and the DM halo spin $\lambda$, the dimensionless form of DM angular momentum. We investigate the connection between halo spin $\lambda$ and galaxy properties in the presence/absence of stellar bars, using the cosmological magneto-hydrodynamic TNG50 simulations at multiple redshifts ($10<z_{r}<$ 1). We determine the bar strength (or bar amplitude, $A_{2}/A_{0}$), using Fourier decomposition of the face-on stellar density distribution. We determine the halo spin for barred and unbarred galaxies ($0.70<A_{2}/A_{0}<$ 0.7) in the centre of the DM halo, close to the galaxy’s stellar disk. At $z_{r}=0$, there is an anti-correlation between halo spin and bar strength. Strongly barred galaxies ($A_{2}/A_{0}>0.4$) reside in DM halos with low spin and low specific angular momentum at their centers. In contrast, unbarred/weakly barred galaxies ($A_{2}/A_{0}<0.2$) exist in halos with higher central spin and higher specific angular momentum. The anti-correlation is due to the barred galaxies’ higher DM mass and lower angular momentum than the unbarred galaxies at $z_{r}=0$, as a result of galaxy evolution. At high redshifts ($z_{r}=1$), all galaxies have higher halo spin compared to those at lower redshifts ($z_{r}=0$), with a weak anti-correlation for galaxies having $A_{2}/A_{0}>0.2$. The formation of DM bars in strongly barred systems highlights how angular momentum transfer to the halo can influence its central spin.
@@ 1 Introduction
Investigating the properties of halo spin and its connection with galaxy properties has been of interest in many studies (Bett et al., 2007, 2010; Bett & Frenk, 2012, 2016). Halo spin plays an important role determining galaxy sizes, as faster-rotating halos having higher halo spins are associated with larger stellar disks (Kim & Lee, 2013; Jiang et al., 2019; Rodriguez-Gomez et al., 2022; Pérez-Montaño et al., 2022; Yang et al., 2023). Romeo et al. (2023) studied a large sample of barred S0 galaxies and found a connection between the specific angular momentum of the disk, galaxy morphology and bar structure. Zjupa & Springel (2017) found that the halo spin (both $\lambda_{B}$ and $\lambda_{P}$ at $r_{vir}$) are not much affected by the collapse of the baryons in the DM halos. Recently, Rosas-Guevara et al. (2022) using current cosmological hydrodynamical TNG50 simulations (Nelson et al., 2019; Pillepich et al., 2019) showed that the halo spin (at $r_{vir}$; $\lambda_{B}$) of barred galaxies is lower than that of the unbarred galaxies at redshift $z_{r}=0$ (also see Izquierdo-Villalba et al. 2022). However, there has been no comprehensive study on the origin of the connection between bar strength and halo spin or halo angular momentum. We will see later in Section 4.1, that the angular momentum of the stellar disk is more than the angular momentum of the DM halo in the central region of the galaxies (for example, 10 times more inside $r<10$ kpc). Only a significant transfer in angular momentum from the disk to the DM halo can lead to a halo spin change in the halo’s inner region. However, no amount of angular momentum transfer can change the halo spin within the virial radius of the halo.
@@ 2 TNG50 Simulations and galaxy sample
We avoid galaxies that do not have enough star particles to resolve the central region of the disk. Following the above criteria, we select barred and unbarred disk galaxies from the above samples, with 597 galaxies at redshift $z_{r}=0$, 650 galaxies at $z_{r}=0.1$ and 509 galaxies at $z_{r}=1.0$. Note that we not only track the progenitors of the galaxies at higher redshifts but also construct independent samples of galaxies at the three redshifts.
@@ 2.1 Barred and unbarred galaxies
Figure 2: The anti-correlation between bar strength and dark matter halo properties at redshift $z_{r}=0.0$. Figure shows the median halo spin (panel a), median DM angular momentum (panel b), median DM mass (panel c) and median specific angular momentum of DM (panel d) for each of the bar strength bins from Table 1 at redshift $z_{r}=0$, for two radii, $r_{m}=5$ kpc (green, Sample 1r5), 10 kpc (magenta, Sample 1r10), in the velocity range $200140<\rm Max\left[V_{c,DM}(r<r_{m})\right]/(kms^{-1})<$ 200. The solid circles denote the median values and the shaded regions are bounded by the $16^{th}$ and $84^{th}$ percentile curves.
@@ 3 Halo spin of Barred and Unbarred galaxies
We aim to study the halo spin of barred and unbarred galaxies having similar masses of their DM halos in the central region and close to the baryonic disk. We use the circular velocity ${\rm V_{c,DM}}(r)=\sqrt{GM_{DM}(r)/r}$ of the DM component to select galaxies with similar DM halo mass profiles ($M_{DM}(r)$). We use the selection criteria of $200140<\rm Max\left[V_{c,DM}(r<r_{m})\right]/(kms^{-1})<$ 200 at radii $r_{m}=5$ and 10 kpc, on each of the seven bar strength bins at $z_{r}=0$ from Table 1, as it maximises the number of galaxies following this selection criterion. At $z_{r}=0$, we find 25, 21, 42, 34, 29, 29, and 10 galaxies at $r_{m}=10$ kpc (Sample 1r10, hereafter) and 19, 22, 48, 44, 45, 33 and 9 galaxies at $r_{m}=5$ kpc (Sample 1r5, hereafter) for each of the bar strength intervals from Table 1. Except for the highest bar strength bin ($0.70.6<A_{2}/A_{0}<$ 0.7), all other bar strength intervals have a moderate number of galaxies.
@@ 3.1 Galaxies with similar DM circular velocities at redshift $z_{r}=0$
In Figure 2, we present the median halo spin $\lambda(r)$ (panel a), median DM angular momentum $L_{DM}(r)=\sqrt{L^{2}_{x}+L^{2}_{y}+L^{2}_{z}}$ (panel b), median DM mass $M_{DM}(r)$ (panel c) and the median DM specific angular momentum $L_{DM}(r)/M_{DM}(r)$ (panel d) using Sample 1r5 (green) and Sample 1r10 (magenta), with their median bar strength in the x-axis of each panel. The solid circles show the median values of each of the above quantities, and shaded regions are bounded by $16^{th}$ and $84^{th}$ percentile curves. Note that, in all of our analysis, we determine the total halo spin at multiple radii using the definition from Equation 1 in the entire article, and not any derived halo spin (for example, $\lambda_{P}$ or $\lambda_{B}$).
@@
Figure 3: Halo spin of strongly barred galaxies ($A_{2}/A_{0}>0.3$) consistently decreases from high to low redshifts, $z_{r}=1$ to 0. Figure shows the evolution of halo spin $\lambda$ (panel a), DM halo mass $M_{DM}$ (panel b), DM halo angular momentum $L_{DM}$ (panel c), DM specific angular momentum $L_{DM}/M_{DM}$ (panel d), stellar to DM mass ratio $M_{\star,90}/M_{DM,90}$ within $R_{\star,90}$ (panel e), stellar mass $M_{\star,90}$ (panel f), stellar angular momentum $L_{\star,90}$ (panel g) and stellar disk specific angular momentum $L_{\star,90}/M_{\star,90}$ (panel h) of the galaxies in Sample 1r10 at $z_{r}=0$ (magenta), 0.2 (green), 0.5 (blue), 0.7 (orange) and 1 (grey). To avoid crowding, we show three redshifts in all panels except in panel a.
@@ 4.2 Correlations between bar strength and stellar and DM angular momentum
Figure 6: Increase in bar strength leads to the decrease in disk angular momentum and coincides with the rise in DM angular momentum. X-axis of all panels: (a) Pearson correlation coefficient $R$ for stellar angular momentum $L_{\star}$ and bar strength $A_{2}/A_{0}$; (b) DM angular momentum $L_{DM}$ and $L_{\star}$; (c) $L_{\star}$ and stellar mass $M_{\star}$; (d) $L_{DM}$ and DM mass $M_{DM}$ and (e) $L_{DM}$ and $A_{2}/A_{0}$. All quantities are estimated within $r=10$ kpc. The y-axes show the number of systems for barred (green) and unbarred galaxies (yellow).
@@ 4.3 Shadow Dark Matter Bar
In Figure 7 panel b, the two-lobed structure in $\Sigma_{\star}^{sub}(r)$ map is due to the asymmetric structure of the stellar bar in the $\Sigma_{\star}(r)$ map (panel a). In panel c, we see a similar over-density in the underlying DM $\Sigma_{DM}^{sub}(r)$ map that is aligned with the stellar bar. This two-lobed structure indicates a bar in the DM density distribution, and we have checked that it is present for all the strongly barred galaxies in Sample 1r10 with $A_{2}/A_{0}>0.6$ at $z_{r}=0$. We follow the evolution of the two-lobed structure in the DM distribution and find that it starts after the bar formation time. For unbarred galaxies, such a structure is not observed (for example, see panel f). The $\Sigma_{DM}^{sub}(r)$ maps are noisy due to the resolution of the simulation. Ash et al. (2024) recently characterized the DM bar strength for the barred galaxy sample from Rosas-Guevara et al. (2022), and found that the pattern speed of the stellar and DM bar in a galaxy in Subbox-0 (a sub-volume in TNG50 with a box size of 7.5 $h^{-1}$ Mpc) is nearly synchronous, and the DM bar is aligned with the stellar bar during 8 Gyrs of evolution. Using high-resolution zoom-in simulations, we will be able to study DM bars with a level of accuracy comparable to that achieved in high-resolution N-body simulations (for example, see Figure 7 in Collier et al. 2019a) in the future.
@@ 5 Halo spin of galaxies in Sample-2 and Sample-3
We investigate if there is an anti-correlation between the halo spin and the bar strength for galaxies with comparable DM masses at high redshifts, similar to $z_{r}=0$ (Section 2.1). Here we examine the barred and unbarred galaxies at higher redshifts $z_{r}=0.1$ and 1.0 by forming samples of galaxies having similar DM mass using Sample 2 and Sample 3 from Table 1. We apply the same selection criteria from Section 3.1 based on the DM circular velocities: $200140<\rm Max\left[V_{c,DM}(r<r_{m})\right]/(kms^{-1})<$ 200, on each of the seven bar strength bins in Sample 2 at $z_{r}=0.1$ and Sample 3 at $z_{r}=1$. We find 37, 58, 46, 39, 40, 28 and 12 galaxies at $r_{m}=10$ kpc (Sample 2r10 hereafter) and 37, 40, 42, 42, 49, 34, 12, galaxies at $r_{m}=5$ kpc (Sample 2r5 hereafter). At $z_{r}=1.0$, we find lesser number of galaxies in Sample-3 in the above circular velocity range: 5, 7, 11, 19, 23, 28 and 11 galaxies for $r_{m}=5$ kpc (Sample 3r5 hereafter) and 8, 19, 15, 24, 33, 38 and 15 galaxies for $r_{m}=10$ kpc (Sample 3r10 hereafter). We present the median halo spin, median DM angular momentum, median DM mass and median DM specific angular momentum for the seven bar strength bins in Figure 8, with Samples 2r5 (panels a–d, green), 2r10 (panels a–d, magenta), 3r5 (panels e–h, green) and 3r10 (panels e–h, magenta), (similar to Figure 2). We have also conducted tests with a slightly different circular velocity selection criterion of $180130<\rm Max\left[V_{c,DM}(r<r_{m})\right]/(kms^{-1})<$ 180 with larger number of galaxies at $z_{r}=1$ to find 18, 24, 16, 39, 33, 32 and 14 galaxies for $r_{m}=10$ kpc and 8, 12, 12, 24, 30, 35 and 15 galaxies for $r_{m}=5$ kpc for the seven bar strength bins and found similar trends in the results.
@@ 6.1 Sample selection and biases
Figure 9: Dark matter halo mass (left column), angular momentum (middle column) and energy (right column) of the sample galaxies at redshifts $z_{r}=$0.0 (top row), 0.1 (middle row) and 1.0 (bottom row), within spherical radius 10 kpc. In each panel, grey points are values of each quantity divided into seven bar strength bins (Table 1). Blue, green and purple lines trace the median values in each bar strength bin, and the shaded regions show the 16th and 84th percentile of the sample. Each bar strength interval consists of DM halos with a wide range of mass, angular momentum and energy, however, the scatter decreases at high redshifts.
@@ 6.2 Convergence
The galaxies that do not satisfy the convergence criteria at $r_{m}=R_{\star,60}$ are mostly the low-mass galaxies that are not included in the galaxy samples in Sections 3.1 and 5, where we use $r_{m}=5$ and 10 kpc and use massive galaxies with $V_{c,DM}(r<r_{m})>140$ km s-1. Low-mass galaxies have small sizes. For low-mass galaxies, $R_{\star,60}$ are smaller compared to the high-mass galaxies (see Figure 13 in Appendix C). We exclude these low-mass galaxies while estimating halo spin for $r_{m}=R_{\star,60}$ at $z_{r}=0.0,$ 0.1 and 1.0 in Figure 14 (panels c, f and i ) in Appendix D. For other values of $r_{m}$ (i.e., 5 kpc, 10 kpc, $R_{\star,90}$, $R_{200}$, $0.15R_{200}$ and $0.05R_{200}$) the low-mass galaxies satisfy the convergence criteria. For more details on how we avoid the effects of numerical relaxation at small radii in our samples, see Appendix C.
@@ Appendix A Halo spin in massive and low-mass galaxies
We study the DM halo spin for comparable DM mass galaxies within the velocity range of $250200<\rm Max\left[V_{c,DM}(r<r_{m})\right]/(kms^{-1})<$ 250 for $r_{m}=5$ and 10 kpc. We find 0, 6, 21, 23, 33, 8 and 0 number of galaxies for $r_{m}=10$ kpc for the seven bar strength bins in Table 1 and similarly for $r_{m}=5$ kpc we have 0, 10, 19, 19, 17, 3 and 0 galaxies for each of the seven bar strength bins. We avoid low statistics by considering the bar strength intervals with $\geq 6$ galaxies. In Figure 10, we present the halo spin (panel a), halo angular momentum (panel b), halo mass (panel c) and halo specific angular momentum (panel d) for the above sample of barred and unbarred galaxies.